A squash ball hits a wall with an initial speed of u and it rebounds in the opposite direction with speed v. Knowing that some energy has been lost through heat in the collision so that v < u and that the duration of the collision is t, find the magnitude of the average force exerted by the wall on the ball.


a) F = m(u+v)/t

b) f = m(u-v)/t

c) F = 0

d) F = m(2v-u)/t


1
Expert's answer
2022-04-11T12:03:50-0400

Let's use second Newton's law "F=ma" ,

where F and a are average values and let's select positive direction along force;

in this coordinates velocity before collision will be negative

"a=\frac{\Delta v}{t}=\frac{v-(-u)}{t}=\frac{u+v}{t}\\
F=\frac{m(u+v)}{t}"

Answer: a)